Soybean meal is 16% protein;cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320-lb mixture that is 15% protein?
Answer by mananth(16946)

Soybean meal is 16% protein;cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320-lb mixture that is 15% protein?
percent----------------quantity
Soyabean 16---------------- x
Corn 8---------------- 320-x
Total 15---------------- 320
16x+8(320-x)=15 *320
16x+2560-8x=4800
16x-8x=4800-2560
8x=2240
/8
x=280 pounds Soyabean
40 pounds corn